OpenStudy (anonymous):

Having trouble with this one as well.. 27^{x}/9^{2x-1} = 3^{x+4} I came up with: 3^{3x}/3^{2(2x-1)} = 3^{x+4} 3^{2(2x-1) -3x} = x+4 4x-2-3x = X+4 x-2 = x+4 and so on...but this is wrong and i'm not sure why. what did I do wrong?

OpenStudy (anonymous):

\[\frac{3^{3x}}{3^{4x-2}} = 3^{x+4}\]

OpenStudy (anonymous):

\[\frac{3^{3x}}{3^{4x}}* 3^2=2^x * 3^4 \]

OpenStudy (anonymous):

\[3^{3x -4x} = 3^x * 3^2\]

OpenStudy (anonymous):

3^(3x)/3^(6x - 3) = 3^(x+4) 3^(3x-(6x-3)) = 3^(x+4) 3x-6x+3 = x + 4 -4x = 1 x = -0.5 look in your second row : you should have subtract the nominator from the denominator also dont remove the basis from the right side before you do it on the left side.

OpenStudy (anonymous):

and its not -0.5 its -0.25 sorry

OpenStudy (anonymous):

it is \(3^x\) not \(2^x\) typo

OpenStudy (anonymous):

Ishaan, how did you write them as rationals, I cannot figure it out..

OpenStudy (anonymous):

on i was wrong again

OpenStudy (anonymous):

3^(3x)/3^(4x - 2) = 3^(x+4) 3^(3x-(4x-2)) = 3^(x+4) 3x-4x+2 = x + 4 -2x = 2 x = -1

OpenStudy (anonymous):

put 3^x in \[ and \ ] (without space )

OpenStudy (anonymous):

actually correct syntax is 3^{x}

OpenStudy (anonymous):

or try it with equation editor